# DT-DDNN: A Physical Layer Security Attack Detector in 5G RF Domain for CAVs

### v2

@@ III System Model and Problem Formulation
III-A Background
The SSB is usually transmitted within the 5G radio frame using 4 OFDM symbols [32] and includes two synchronization signals: the primary synchronization signal (PSS) and secondary synchronization signal (SSS) [33]. The SSB comprises information necessary for the UE to establish synchronization with the cell, such as the physical cell identity (PCI) of the cell, and additional information provided by physical broadcast channel (PBCH) and physical downlink shared channel (PDSCH)[33]. PSS and SSS use the same time slots as the PBCH. PBCH symbols are concentrated in two or four slots, which gives the appearance of a low-duty cycle, particularly at higher subcarrier spacing. The Master Information BLock (MIB), containing critical data transported by the PBCH, comprises parameters that are vital for the UE to establish a connection with a cell. When a UE powers on or enters a new cell’s coverage area, it uses this information to discover the available SSBs. Scanning the predefined SSB locations within the radio frame, the UE executes an SSB discovery procedure to accomplish this. The UE is capable of decoding the information carried by an SSB, including the PCI, once it has been detected.
@@ III-B Problem formulation
| Output of DWT stage | 11, 10, 17 | |
| Energy per null resource elements, | 14, 15, 16, 17 | |
| log transform of energy, vectorization | ||
| of log transform of energy, respectively | ||
| 3-D tensor of observation | 17, 19, 20, | |
| 21, 22 | ||
| Corresponding label for observation | 18, 21, 22 | |
| Score of the jamming detection algorithm | 19, 20, 21, | |
| for class | 22, 24, 25 | |
| DNN model threshold | 26, 27 |
In the first step of gaining access to a gNB, UE requires information called System Information Block 1 (SIB1) which depends on the decoding of Master Information Block (MIB) [34]. This procedure is achievable only through the detection of SSB. One or more SSBs are transmitted through an SS burst periodically based on a pre-determined periodicity at a five-millisecond window [35]. Each SSB contains cell ID information which is calculated by . Where represents group ID and is related to sector ID [8]. Sector ID is provided by PSS to help reach a coarse time and frequency synchronization. A 5G frame includes numerous slots, each being divided into a specific number of symbols which depends on the subcarrier spacing and numerology. SSB is transmitted through four OFDM symbols in the time domain and spans over subcarriers in the frequency domain (Fig. 1). 127 subcarriers in the first symbol are dedicated to PSS, and there are 113 unused subcarriers below and above PSS which are set to ’’. PSS follows on of the three base -symbols m-sequences in frequency domain. Each m-sequence is a circular shift version of the other two and their cross-correlation value is equal to zero [36]. The three base m-sequences are demonstrated in (1).
@@
In the scenario where a jamming signal () is introduced, the received signal becomes susceptible to the impact of the jammer. Detection of the jamming signal can be represented as a binary hypothesis framework (7),(8), where we have a null hypothesis denoted as and an alternative hypothesis, , corresponding to the presence of the jamming signal.
@@ IV-A1 PSS correlation
Each transmitted SSB includes a PSS sequence which is a version of one of the three base PSS sequences formulated in (1). Thus, within the correlation signal of the received PSS sequence with one of the m-sequences, we expect to observe a visible peak value. The correlation of with each m-sequences results in a 1 by 2 signal demonstrated as444Since there is a specific signal for each , the output of the correlator in (9) is sub-scripted with . ,
@@ IV-A2 DWT
where is derived as per (9). Furthermore, is the mother wavelet function chosen as Haar wavelet 555Haar wavelet is a set of rescaled square-shaped functions that collectively form the simplest wavelet family or basis [37]. to maintain and magnify the important information from PSS correlation. Haar wavelet is a straightforward choice as it is very effective at identifying abrupt changes in the signal amplitude including PSS peaks [38][39]. Thus, it can successfully reduce signal size while maintaining crucial information and keeping the computational load at a desired stage. Fig. 3 demonstrates the effect of adding two layers of DWT to the PSS correlation dataset. As can be seen in the figure, the length of the correlated signal is decreased by almost 25%, and the correspondent peak is magnified significantly by 2 times.
IV-A3 Energy per Null RE (EPNRE)
@@
in which represents the Fourier transform of the received signal in -th OFDM symbol and -th subcarrier data point. In the next step, the SS block is extracted from the frequency domain signal. There are numbers of resource elements in the SS block which are intentionally set to zero by gNB as shown in Fig. 1. These resource elements help us collect information about the noise and jamming signal. The energy of these resource elements is calculated as expressed in (14).
Depending on the jammer transmit power and its distance to the UE, the received energy of the jamming signal might be small in some observations. Thus, to adjust the dynamic range of the calculated received energy to be more distinguishable, a log function is applied to the final energy value as,
@@ IV-B1 DNN Input
Observations are designed as a 2-D image including the PSS correlation signals at the output of the second DWT module (), and energy per null resource elements (). In each observation (), the first three rows in the 2-D image include correlated signal, and the last two rows are dedicated to the energy parameter (Fig. 2). The calculated EPNRE is a scalar value while the size of is equal to . Thus, to maintain the impact of the energy feature and to make it visible to the classifier, the energy is repeated in the last two rows as,
where denotes a vector of ones. After creating a 2-D image of features, each observation is positioned behind the previous observation thus creating a 3-D tensor (Fig. 2) as the final dataset ().
@@ IV-B2 Data Augmentation and Class Balancing
The dissimilarity in sample sizes between the two classes, although frequently seen in the process of data collection, causes the class imbalance that must be resolved to maintain the reliability and validity of the produced data. These imbalanced classes can result in biased models and fail to generalize to minority class [40][41]. This situation causes the CNN model to demonstrate less efficiency and accuracy in classifying the minority observations and decreases the performance of the model in detecting underrepresented received signals. Various techniques, including oversampling the training samples to produce a comparatively balanced dataset [42], have been suggested as solutions to this problem. Additionally, research has demonstrated that CNN decisions are more difficult to generalize to minority classes when the data is unbalanced [43]. However, methods including data augmentation have demonstrated efficacy in enhancing the precision of classification tasks involving unbalanced datasets [44], [45], [46], [47], [48]. while adversarial machine learning approaches have shown promising solutions in securing wireless systems through augmentation of real data samples [49], qualitative and quantitative analyses of various data augmentation techniques are provided in several works, notably in [50], [51], [52], [53]. To overcome the effect of the imbalance classes and use a more robust dataset, circular shift is applied to the minor dataset [54]. Through the circular shift augmentation method, each observation in the minority class is divided into sub-sequences with similar lengths, and each sub-sequence is independently shuffled to create a new sequence which introduces new variations of the collected observations. Traditional techniques (such as rotation, cropping, and flipping) may affect data integrity or eliminate vital information. Circular shift enhances the generalization capabilities of the DNN model by introducing diverse patterns while maintaining most of the original information without significant information loss [54], [55],.
@@ IV-C2 Deep Cascade Learning
Although the convolution network has demonstrated remarkable performance and results [56] [57][58][59], it experiences the vanishing gradient problem while undergoing training [60]. This occurs because the weight updates during back-propagation are substantially reduced as the depth of the network increases, meaning that layers closer to input experience a slower learning rate [61],[62]. One solution to overcome this situation is the deep cascade learning algorithm proposed in [61]. Deep cascade learning gradually trains the network from the lowest to higher layers. When the jamming signal is weak and therefore SJNR values are close to each other, the classifier faces trouble classifying between two classes. Deep cascade learning fine-tunes the weights by dividing the network into sub-layers and sequentially trains each layer until all the input layers have been trained. Through this technique, the vanishing gradient problem can be eliminated by compelling each layer of the network to acquire features that are correlated with the output. In other words, it maintains the linear correlation between input and output while accommodating the nonlinear relationship [63]. Furthermore, it has shown a significant reduction in training time and memory while adjusting the complexity of the network to the given data [61][64][65]. The structure of the DNN model using deep cascade learning is shown in Fig. 6. The input layer is connected to the output using two dense layers. The weights for the initial model layer and the output are subsequently obtained via a back-propagation algorithm during the training. After reaching a state of stability, the second layer is trained by connecting to an output layer with a similar structure to the first layer using data created by forward propagating the actual inputs through the fixed initial layer. Through several iterations, every layer goes through this process for learning, and the weights remain constant for the subsequent layer. Adapting this approach helps increase the robustness of the second DNN and preserve the back-propagated gradient while having hidden layers.
V Simulation and Experiment Results
In this section, dataset generation and the performance of the proposed deep jamming detection architecture in offline training and online testing are evaluated and discussed thoroughly.
@@ V-A1 Data generation
To collect the received signal recordings, a synthetic 5G dataset is generated from the scenarios with and without jammer under different conditions. The generated dataset is then processed by data preprocessing block and fed into the detection model. A detailed description of the dataset generation environment is provided as follows.
@@
where and signify the wavelength and the distance between gNB and UE, respectively. The received signal for each observation is recorded in time domain and the number of observations is . The same process is performed to record the received signal in the presence of the jammer with different SJNR values and jamming powers. The number of observations in this scenario is denoted as . The jammer used for the training is designed using AWGN. Other types of jammers such as the jammer which can transmit modulated BPSK, and 8QAM signals are used for testing. The dataset parameters used for the dataset generation are extracted from [66], [67] and listed in Table IV.
@@
| Sample Rate666The sample rate equals to which equals Msps | 61.44 Msps |
@@ V-A2 Performance under the Single Threshold DNN
Fig. 8 presents the confusion matrix for jamming detection with a single threshold DNN design while the SJNR is spanning from -10.5 to 30 dB. The figure demonstrates the performance of the detection algorithm in which true positive (TP) and true negative (TN) correspond to non-jammed and jammed classes respectively. Based on this matrix, the algorithm classified 5322 observations as TP with the rate of 88.0%, and 5201 cases as TN with the rate of 86.0%. 12.0% of the jammed cases were falsely classified as non-jammed scenarios and 14.0% of the non-jammed observations were erroneously classified as jammed cases. This demonstrates the ability of the model to accurately differentiate between jammed and non-jammed signals. However, it also introduces the potential for further improvement to decrease the misclassification cases particularly to reduce the FN observations. In the jamming detection scenario, reducing FN is more important than FP, as failure to identify a jammer leads to security risks and communication disruption [68].
After training and testing the model, a different test dataset is implemented consisting of signals transmitted over CDL-C and CDL-A channel models with delay spread. The purpose of this test is to show the generalization of the proposed model to various 5G network settings. The confusion matrix in Fig. 9 demonstrates the robustness of the design to the disparities of the network by correctly detecting 97.3% of the non-jammed cases and 94.1% of jammed cases. Furthermore, recall and precision can be obtained as 0.97 and 0.94 respectively. Hence, the model demonstrates minimal misclassification which shows its ability to detect the attack despite increased complexity. This test highlights the potential of the detection model in a real-world 5G network while maintaining performance and accuracy in a more complex test environment.
Besides testing under different network settings, a second test dataset compromised by a jammer transmitting with 8QAM signal modulation is created to test the effectiveness of the detector model to different types of jamming signals. Based on the confusion matrix of this test which is provided in Fig. 10, despite using AWGN as the jamming signal in the training process, it demonstrates acceptable flexibility by accurately classifying 97% of non-jammed cases and 100% of jammed observations. These results show the robustness of the design against novel jamming techniques which signifies a transfer of learning from one jamming mode to another. However, no percentage of jammed cases were incorrectly classified, which corresponds to zero probability of miss-detection (FP=0). These results demonstrate the ability of the model to stay generalized to various types of jamming without requiring retraining.
@@ V-A3 DT-DDNN Performance
The confusion matrix of the DT-DDNN design which includes a double DNN architecture to improve the sensitivity of the jamming detector is presented in Fig. 12. While the first DNN uses a double-threshold concept to increase the sensitivity, the second DNN includes a deep cascade learning model to enhance the ability of the system to accurately classify complex observations and decrease the probability of miss-detection to 9.0%. The model accurately classifies 91.0% of non-jammed cases and 96.4% of jammed observations, which shows almost 3% and 10.4% rise in the accuracy compared to the single threshold architecture. These results represent an advancement in the ability to detect and provide countermeasures for jamming attacks in real-world scenarios of the 5G network.
@@
Comparative analysis of three DNN designs for jamming detection in a 5G network is provided by the ROC curves in Fig. 13. The green dotted curve corresponds to a DNN trained on raw IQ samples taken from 5G waveform with no preprocessing block in the design. The optimum configuration was achieved with 4 layers of CNN with batch and ReLu following each layer. The single threshold DNN design, illustrated by the blue dashed curve, is correspondence to the single threshold DNN design that uses a single DNN to classify jammed cases. In contrast, the red solid curve represents the DT-DDNN design which uses a double DNN system to improve the sensitivity of the detection algorithm. The primary DNN block detects jamming by applying a double threshold decision-making method, and the secondary DNN handles observations that fall into the ambiguous area that the first DNN has found difficulties in classifying. By integrating a deep cascade learning model into the second DNN, the objective is to further enhance the classification performance. Based on the ROC curves, it is evident that the DT-DDNN provides a higher probability of detection () in comparison to the majority of false alarm probabilities (). This indicates a greater proportion of true positive and a reduced potential of miss-detection of legit signals as jamming. This suggests that the second design with DT-DDNN architecture offers a more robust jamming detection system under challenging detection cases.

### v3

@@ III System Model and Problem Formulation
III-A Background and Feasibility Analysis
The SSB, illustrated in Fig. 1(a), is usually transmitted within the 5G radio frame using 4 OFDM symbols [34] and includes two synchronization signals: the PSS and SSS [35]. The SSB comprises information necessary for the UE to establish synchronization with the cell, such as the PCI of the cell, and additional information provided by the PBCH and physical downlink shared channel (PDSCH)[35]. PSS and SSS use the same time slots as the PBCH. PBCH symbols are concentrated in two or four slots, which gives the appearance of a low-duty cycle, particularly at higher subcarrier spacing. The MIB, containing critical data transported by the PBCH, comprises parameters that are vital for the UE to establish a connection with a cell. When a UE powers on or enters a new cell’s coverage area, it uses this information to discover the available SSBs. Scanning the predefined SSB locations within the radio frame, the UE executes an SSB discovery procedure to accomplish this. The UE is capable of decoding the information carried by an SSB, including the PCI, once it has been detected. This detection process is performed by analyzing the correlation of the received SSS and PSS signals and the known base SSS and PSS signals. A jammer that impacts the SSB of 5G communication can disrupt the PCI detection and synchronization process by a simple interference signal. PSS signals will be described and mathematically modelled in section IV-A1. However, to better understand the effect of the jamming attack on the synchronization process, we provide the output of PCI detection based on PSS, i.e. the cross correlation between received PSS and basis PSS sequences , under different jamming powers in Fig. 1(b). When the SJNR value is equal to , the output of the correlation of received PSS and base PSS sequence results in a clear peak which leads the UE to detect the PCI of the available gNB. As the SJNR decreases and reaches with higher jamming power, the correlation peak detection becomes more challenging, disrupting legitimate communication. In addition, based on the SSB in Fig. 1(a), it can be concluded that some resource elements are set to zero based on the 3GPP standard containing zero energy. When a jammer starts sending jamming signals, the energy of these resource elements starts increasing. This is another characteristic that can help to detect jamming attacks as detailed under the Energy per Null RE (EPNRE) description in section IV-A3. Fig. 1(c) demonstrates the total energy of all null resource elements when the power of jammer decreases and SJNR is increased from to . It is observed that as the jamming power is decreased, the EPNRE decreases toward zero. This change in EPNRE can be used to detect the presence of a jammer in the network.
@@ III-B Problem formulation
| Output of DWT stage | 11, 10, 15 | |
| Energy per null resource elements, | 14, 15 | |
| vectorization of log transform of energy | ||
| 3-D tensor of observation | 15, 17, 18, | |
| 19, 20 | ||
| Corresponding label for observation | 16, 19, 20 | |
| Score of the jamming detection algorithm | 17, 18, 19, | |
| for class | 20, 22, 23 | |
| DNN model threshold | 24, 25 |
In the first step of gaining access to a gNB, UE requires information called System Information Block 1 (SIB1) which depends on the decoding of MIB [36]. This procedure is achievable only through the detection of SSB. One or more SSBs are transmitted through an SS burst periodically based on a pre-determined periodicity at a five-millisecond window [37]. Each SSB contains cell ID information which is calculated by . Where represents group ID and is related to sector ID [8]. Sector ID is provided by PSS to help reach a coarse time and frequency synchronization. A 5G frame includes numerous slots, each being divided into a specific number of symbols which depends on the subcarrier spacing and numerology. SSB is transmitted through four OFDM symbols in the time domain and spans over subcarriers in the frequency domain. 127 subcarriers in the first symbol are dedicated to PSS, and there are 113 unused subcarriers below and above PSS which are set to ’’. PSS follows on of the three base -symbols m-sequences in frequency domain. Each m-sequence is a circular shift version of the other two and their cross-correlation value is equal to zero [38]. The three base m-sequences are demonstrated in (1).
@@
In the scenario where a jamming signal () 444As the focus of our detection system is investigating the impact of the jammer on 5G SSB, the jamming signal is modeled based on the AWGN noise that reflects both smart SSB jammer and barrage jammer. is introduced, the received signal becomes susceptible to the impact of the jammer. Detection of the jamming signal can be represented as a binary hypothesis framework (7),(8), where we have a null hypothesis denoted as and an alternative hypothesis, , corresponding to the presence of the jamming signal.
@@ IV-A Data Preprocessing
Preprocessing of data in ML techniques is a vital part. If done appropriately, it helps transfer raw data into a neat feature space that can be exploited by the DL module for high-performance classification. Data preprocessing in the proposed architecture includes three main steps, performing PSS correlation, calculating DWT, and extracting EPNRE. The first step helps detect low-power jamming since the peak amplitude and the correlation output pattern can reflect the presence of the jammer even in small jamming powers as depicted in Fig. 1(b). The DWT acts as a feature space dimension reduction techniques, making the training process more easy and less time-consuming. Extracting EPNRE helps the DL decision be robust to channel behavior as no signal energy from the gNB is assumed to be there.
@@ IV-A1 PSS correlation
Each transmitted SSB includes a PSS sequence which is a version of one of the three base PSS sequences formulated in (1). Thus, within the correlation signal of the received PSS sequence with one of the m-sequences, we expect to observe a visible peak value. The correlation of with each m-sequences results in a 1 by 2 signal demonstrated as555Since there is a specific signal for each , the output of the correlator in (9) is sub-scripted with . ,
@@ IV-A2 DWT
where is derived as per (9). Furthermore, is the mother wavelet function chosen as Haar wavelet 666Haar wavelet is a set of rescaled square-shaped functions that collectively form the simplest wavelet family or basis [39]. to maintain and magnify the important information from PSS correlation. Haar wavelet is a straightforward choice as it is very effective at identifying abrupt changes in the signal amplitude including PSS peaks [40][41]. Thus, it can successfully reduce signal size while maintaining crucial information and keeping the computational load at the desired stage. Fig. 3 demonstrates the effect of adding two layers of DWT to the PSS correlation dataset. As can be seen in the figure, the length of the correlated signal is decreased by almost 25%, and the correspondent peak is magnified significantly by 2 times.
IV-A3 EPNRE
@@
in which represents the Fourier transform of the received signal in -th OFDM symbol and -th subcarrier data point. In the next step, the SS block is extracted from the frequency domain signal. There are numbers of resource elements in the SS block which are intentionally set to zero by gNB [38]. These resource elements help us collect information about the noise and jamming signal. The energy of these resource elements is calculated as expressed in (14).
Depending on the jammer transmit power and its distance to the UE, the received energy of the jamming signal might be small in some observations. Thus, to adjust the dynamic range of the calculated received energy to be more distinguishable by the classifier, a log transform is applied to the final energy value, i.e., .
@@ IV-B1 DNN Input
Observations are designed as a 2-D image including the PSS correlation signals at the output of the second DWT module (), and energy per null resource elements (). In each observation (), the first three rows in the 2-D image include correlated signal, and the last two rows are dedicated to the energy parameter (Fig. 2). The calculated EPNRE is a scalar value while the size of is equal to . Thus, to maintain the impact of the energy feature and to make it visible to the classifier, the energy is repeated in the last two rows. Mathematically speaking, where denotes a vector of ones. After creating a 2-D image of features, each observation is positioned behind the previous observation thus creating a 3-D tensor (Fig. 2) as the final dataset ().
@@ IV-B2 Data Augmentation and Class Balancing
The dissimilarity in sample sizes between the two classes causes the class imbalance that can result in biased models and fail to generalize to minority class [42]. Data augmentation techniques provided in several works, notably in [43, 44, 45] have demonstrated efficacy in enhancing the precision of classification tasks involving unbalanced datasets [46]. Circular shift is one of the data augmentation methods applied to the minor dataset to overcome the effect of the imbalance classes and is introduced and described in detail in [47]. Another reason for performing circular shift is that the PSS correlation peak in (9) might not happen exactly at the center sample of due to imperfect synchronization in time domain and sample rate mismatch between gNB and the receiver. Circular shift helps the DNN model to learn different patterns existed in the real world data. Through the circular shift augmentation method, each observation in the minority class is divided into sub-sequences with similar lengths, and each sub-sequence is independently shuffled to create a new sequence which introduces new variations of the collected observations. Traditional techniques (such as rotation, cropping, and flipping) may affect data integrity or eliminate vital information. Circular shift enhances the generalization capabilities of the DNN model by introducing diverse patterns while maintaining most of the original information without significant information loss [47], [48].
@@ IV-C2 Deep Cascade Learning
Although the convolution network has demonstrated remarkable performance and results [49][50], it experiences the vanishing gradient problem while undergoing training [51]. This occurs because the weight updates during back-propagation are substantially reduced as the depth of the network increases, meaning that layers closer to input experience a slower learning rate [52],[53]. One solution to overcome this situation is the deep cascade learning algorithm proposed in [52]. Deep cascade learning gradually trains the network from the lowest to higher layers. When the jamming signal is weak and therefore SJNR values are close to each other, the classifier faces trouble classifying between two classes. Deep cascade learning fine-tunes the weights by dividing the network into sub-layers and sequentially trains each layer until all the input layers have been trained. Through this technique, the vanishing gradient problem can be eliminated by compelling each layer of the network to acquire features that are correlated with the output. In other words, it maintains the linear correlation between input and output while accommodating the nonlinear relationship [54]. Furthermore, it has shown a significant reduction in training time and memory while adjusting the complexity of the network to the given data [52][55]. The structure of the DNN model using deep cascade learning is shown in Fig. 6. The input layer is connected to the output using two dense layers. The weights for the initial model layer and the output are subsequently obtained via a back-propagation algorithm during the training. After reaching a state of stability, the second layer is trained by connecting to an output layer with a similar structure to the first layer using data created by forward propagating the actual inputs through the fixed initial layer. Through several iterations, every layer goes through this process for learning, and the weights remain constant for the subsequent layer. Adapting this approach helps increase the robustness of the second DNN and preserve the back-propagated gradient while having hidden layers.
V Simulation and Experimental Results
@@ V-A1 Data generation
To collect the received signal recordings, a synthetic 5G dataset is generated using MATLAB R2024a from the scenarios with and without a jammer under different conditions. The generated dataset is then processed by the data preprocessing block and fed into the detection model. A detailed description of the dataset generation environment is provided as follows.
@@
where and signify the wavelength and the distance between gNB and UE, respectively. The received signal for each observation is recorded in the time domain, and the number of observations is . The same process is performed to record the received signal in the presence of the jammer with different SJNR values and jamming powers. The number of observations in this scenario is denoted as . The jammer used for the training is designed using AWGN. Other types of jammers, such as the jammer which can transmit modulated BPSK, and 8QAM signals, are used for testing. The dataset parameters used for the dataset generation are extracted from [56], [57] and listed in Table IV. A total of observations are generated for each class of unjammed and jammed received signals using the parameters listed in Table IV. These observations are evenly distributed across all modulation types and and three PSS indices of .
@@
| Sample Rate777The sample rate equals to which equals Msps | 61.44 Msps |
@@ V-A2 Performance under the Single Threshold DNN
Fig. 8(a) presents the confusion matrix for jamming detection with a single threshold DNN design while the SJNR spans from -10.5 to 30 dB. The figure demonstrates the performance of the detection algorithm in which true positive (TP) and true negative (TN) correspond to non-jammed and jammed classes respectively. Based on this matrix, the algorithm classified 5322 observations as TP with the rate of 88.0%, and 5201 cases as TN with the rate of 86.0%. 12.0% of the jammed cases were falsely classified as non-jammed scenarios and 14.0% of the non-jammed observations were erroneously classified as jammed cases. This demonstrates the ability of the model to accurately differentiate between jammed and non-jammed signals. However, it also introduces the potential for further improvement to decrease the misclassification cases particularly to reduce the FN observations. In the jamming detection scenario, reducing FN is more important than FP, as failure to identify a jammer leads to security risks and communication disruption [58].
After training and testing the model, a different test dataset is implemented consisting of signals transmitted over CDL-C and CDL-A channel models with delay spread. The purpose of this test is to show the generalization of the proposed model to various 5G network settings. The confusion matrix in Fig. 8(b) demonstrates the robustness of the design to the disparities of the network by correctly detecting 97.3% of the non-jammed cases and 94.1% of jammed cases. Furthermore, recall and precision can be obtained as 0.97 and 0.94 respectively. Hence, the model demonstrates minimal misclassification which shows its ability to detect the attack despite increased complexity. This test highlights the potential of the detection model in a real-world 5G network while maintaining performance and accuracy in a more complex test environment.
Besides testing under different network settings, a second test dataset compromised by a jammer transmitting with 8QAM signal modulation is created to test the effectiveness of the detector model to different types of jamming signals. Based on the confusion matrix of this test which is provided in Fig. 8(c), despite using AWGN as the jamming signal in the training process, it demonstrates acceptable flexibility by accurately classifying 97% of non-jammed cases and 100% of jammed observations. These results show the robustness of the design against novel jamming techniques which signifies a transfer of learning from one jamming mode to another. However, no percentage of jammed cases were incorrectly classified, which corresponds to zero probability of miss-detection (FP=0). These results demonstrate the ability of the model to stay generalized to various types of jamming without requiring retraining.
@@ V-A3 DT-DDNN Performance
The confusion matrix of the DT-DDNN design which includes a double DNN architecture to improve the sensitivity of the jamming detector is presented in Fig. 8(d). While the first DNN uses a double-threshold concept to increase the sensitivity, the second DNN includes a deep cascade learning model to enhance the ability of the system to accurately classify complex observations and decrease the probability of miss-detection to 9.0%. The model accurately classifies 91.0% of non-jammed cases and 96.4% of jammed observations, which shows almost 3% and 10.4% rise in the accuracy compared to the single threshold architecture. These results represent an advancement in the ability to detect and provide countermeasures for jamming attacks in real-world scenarios of the 5G network.
The proposed method is compared to various state-of-the-art approaches in the literature. We tailored the proposed methods in [23] and [31] according to our scenario. Similar to [31], we feed unprocessed data (I-Q samples) to the pre-trained CNN-based classifier. Instead of using AE trained with single class, we use CNN-based supervised learning as higher performance was observed in low jamming power. The work in [23] uses Deep Attention recognition mechanism (DatR) with an LSTM network for both LoS and NLoS cases with the input features of SINR and RSSI fed to the ML module. To improve the detection performance in NLoS where the received signal power is low, augmentation and majority voting are combined. Table V compares accuracy, TP, and TN values for single-threshold, DT-DDNN, unprocessed I-Q [31], and DAtR [23]. For the DAtR model, the 1-D structure described in [23] is used, with two window sizes, and , evaluated for comparison. Under test data with high SJNR, the DT-DDNN exhibits the accuracy, TP, and TN of 0.9368, 0.9144, and 0.9617, respectively, representing the highest performance. DAtR with and the proposed single threshold experience similar accuracy. While TP of the former is higher than the latter, DAtR yields lower TN than the single threshold method.
| Method | Accuracy | TP | TN |
| Single Threshold | 0.879 | 0.8775 | 0.8627 |
| DT-DDNN | 0.9368 | 0.9144 | 0.9617 |
| Unprocessed I-Q [31] | 0.8393 | 0.8501 | 0.8300 |
| DAtR [23] | 0.8540 | 0.8700 | 0.8403 |
| DAtR [23] | 0.8719 | 0.9001 | 0.8515 |
@@
Comparative analysis of three DNN designs for jamming detection in a 5G network is provided by the ROC curves in Fig. 10. The green dotted curve corresponds to a DNN trained on raw IQ samples taken from 5G waveform with no preprocessing block in the design. The optimum configuration was achieved with 4 layers of CNN with batch and ReLu following each layer. The single threshold DNN design, illustrated by the blue dashed curve, is correspondence to the single threshold DNN design that uses a single DNN to classify jammed cases. The ROC for DAtR is differentiated with dashed-line in purple in which is used due to improved performance. In contrast, the red solid curve represents the DT-DDNN design which uses a double DNN system to improve the sensitivity of the detection algorithm. The primary DNN block detects jamming by applying a double threshold decision-making method, and the secondary DNN handles observations that fall into the ambiguous area that the first DNN has found difficulties in classifying. By integrating a deep cascade learning model into the second DNN, the objective is to further enhance the classification performance. Based on the ROC curves, it is evident that the DT-DDNN provides a higher probability of detection () in comparison to the majority of false alarm probabilities (). This indicates a greater proportion of true positive and a reduced potential of miss-detection of legit signals as jamming. This suggests that the second design with DT-DDNN architecture offers a more robust jamming detection system under challenging detection cases. It is worth noting that since the received signal is affected by the communication channel, providing an ML module with at least one feature independent of the channel behavior, EPNRE herein, helps improve the detection performance.